# Nontriviality of rings of integral-valued polynomials

### v1

@@ 4 Global Conditions and Examples
Recall that when $\alpha\in\overline{\mathbb{Z}}$, the index of $\alpha$ is $\iota_{\alpha}=[O_{\mathbb{Q}(\alpha)}:\mathbb{Z}[\alpha]]$. The lemma below summarizes the basic relationships among $\Psi_{p,n}$, $\iota_{\alpha}$, and integer-valued polynomials.
@@
    Modulo $p$, $\Psi_{p,n}$ is the monic least common multiple of all polynomials in $\mathbb{F}_{p}[X]$ of degree at most $n$.
@@
Part (1) is shown in [2, Theorem 3]. For (2), if $\alpha\in\mathcal{A}_{n}$ with minimal polynomial $m_{\alpha}(X)$, then $m_{\alpha}\mid\Psi_{p,n}$ modulo $p$. Hence, $\Psi_{p,n}(\alpha)\in pO_{\mathbb{Q}(\alpha)}$. To show that $\Psi_{p,n}/p\notin\textnormal{Int}_{\mathbb{Q}}(\mathcal{A}_{n+1})$, let $f\in\mathbb{F}_{p}[X]$ be monic and irreducible of degree $n+1$. Then, $f\nmid\Psi_{p,n}$ mod $p$. Next, let $F\in\mathbb{Z}[X]$ be monic, irreducible, and such that $F\equiv f$ mod $p$ (the existence of $F$ follows from Perron’s Criterion for irreducibility [27, Theorem 2.2.5]). Let $\beta\in\overline{\mathbb{Z}}$ be a root of $F$, and let $O=O_{\mathbb{Q}(\beta)}$; note that $\beta\in\mathcal{A}_{n+1}$. Suppose that $\Psi_{p,n}(\beta)/p\in\overline{\mathbb{Z}}$. Then, $\Psi_{p,n}(\beta)\in pO$, and hence $\Psi_{p,n}(\beta)\equiv 0$ mod $p$. Since the image of $\beta$ in $O/pO$ is a root of $f$, this means that $f\mid\Psi_{p,n}$ mod $p$, which is a contradiction.
@@
    Let $\sigma\in G=\text{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and $U$ be a valuation of $\overline{\mathbb{Q}}$ extending $\mathbb{Z}_{(p)}$ for some $p\in\mathbb{P}$. Then $\textnormal{Int}_{\mathbb{Q}}(S,U)=\textnormal{Int}_{\mathbb{Q}}(\sigma(S),\sigma(U))$. In particular, if $S$ is $G$-invariant (i.e., $\sigma(S)=S,\forall\sigma\in G$) and $U,U^{\prime}$ are two valuation domains extending $\mathbb{Z}_{(p)}$ then $\textnormal{Int}_{\mathbb{Q}}(S,U)=\textnormal{Int}_{\mathbb{Q}}(S,U^{\prime})$ and so $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})=\textnormal{Int}_{\mathbb{Q}}(S,U)$ for any valuation domain $U$ extending $\mathbb{Z}_{(p)}$.
@@
    Assume that for each integer $n\geq 1$, there exists a finite subset $\{s_{1},\ldots,s_{m}\}\subseteq S$ such that $[\mathbb{Q}(s_{i}):\mathbb{Q}]>n$ for each $1\leq i\leq m$ and $\gcd(\iota_{s_{1}},\ldots,\iota_{s_{m}})=1$. Then, $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ is trivial.
@@
Part (1) is Example 1.3. For (2), let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ of degree $d$ and let $s_{1},\ldots,s_{m}$ be the elements of $S$ with the stated properties. For each $i$, we have $[\mathbb{Q}(s_{i}):\mathbb{Q}]>d$, so $\iota_{s_{i}}f\in\mathbb{Z}[X]$ by Lemma 4.2(3). The condition $\gcd(\iota_{s_{1}},\ldots,\iota_{s_{m}})=1$ then implies that $f\in\mathbb{Z}[X]$. ∎
@@
    If $m\in\mathbb{N}$ and $q$ is a prime such that $q\nmid c$, then there exists a prime $l>m$ such that $q\nmid\iota_{c,l}$.
@@
(1) This is true by from Perron’s Criterion for irreducibility [27, Theorem 2.2.5].
@@
(3) Assume $p\mid c$, and use a bar to denote reduction mod $p$. Then, $\overline{f_{c,n}}(X)=X^{n}$, and $f_{c,n}(X)=X^{n}+pF_{c,n}(X)$, where $F_{c,n}(X)=\tfrac{c^{3}}{p}X^{n-1}+\tfrac{c^{2}}{p}$. Note that $F_{c,n}\in\mathbb{Z}[x]$ and $\overline{F}=0$. By Theorem 4.8, $p\mid\iota_{p,n}$.
@@
(5) We have $f_{c,n}^{\prime}(X)=X^{n-2}((n-1)c^{3}+nX)$. By the Dirichlet’s theorem on arithmetic progressions, there exists a prime $l$ of the form $qk+1$, $l>m$; in particular, $v_{q}(l-1)>0$ and so, modulo $q$, the only root of $f_{c,l}^{\prime}(X)$ is 00 (note that $v_{q}(l)=0$) which is not a root of $f_{c,l}$ modulo $q$ because we are assuming $v_{q}(c)=0$. Hence, $f_{c,l}$ is separable modulo $q$ so that $q\nmid\iota_{c,l}$ by Theorem 4.8. ∎
@@
By Lemma 4.10(2), $u_{p}(s_{i})=2/q_{i}$ for each $i$, so $\{u_{p}(s_{i})\}_{i\in\mathbb{N}}$ is a strictly decreasing sequence with limit 00. By Lemma 4.6, $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})=\mathbb{Z}_{(p)}[X]$.
Let now $q$ be a prime different from $p$. Given $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$, by Lemma 4.10(5), there exists a prime $q_{i}$ such that $[\mathbb{Q}(s_{i}):\mathbb{Q}]=q_{i}>\deg(f)$ and $q\nmid\iota_{s_{i}}$. By Lemma 4.2(3), $\iota_{s_{i}}\cdot f\in\mathbb{Z}[X]$ so that $f$ belongs to $\mathbb{Z}_{(q)}[X]$. This argument shows that $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(q)}})=\mathbb{Z}_{(q)}[X]$ for every prime $q\not=p$.
@@
We recall that the prime $p$ is totally ramified in $\mathbb{Q}(\zeta_{p^{k}})$ for every $k\in\mathbb{N}$ (see for example [16, Theorem 26]), and $u_{p}(1-\zeta_{p^{k}})=(p^{k-1}(p-1))^{-1}$ by [29, Chapter 2, p. 9]. Moreover, if $q$ is a prime different from $p$ then $u_{q}(1-\zeta_{p^{k}})=0$. These calculations prove the claims about $S$.
@@
Given primes $p<q$, we have $\zeta_{q}-\zeta_{p}=\zeta_{q}(1-\zeta_{pq}^{q-p})$, and $\zeta_{pq}^{q-p}$ is a primitive $pq$-th root of unity because $q-p$ is coprime to $pq$. So, for any prime $l$,
@@
By [29, Proposition 2.8], $1-\zeta_{pq}$ is a unit of $\mathbb{Z}[\zeta_{pq}]$. Hence, $u_{l}(\zeta_{q}-\zeta_{p})=0$, and so for every prime $l$, $S$ is pseudo-stationary with respect to $u_{l}$ with breadth ideal $U_{l}$.
@@
Let =$\mathbb{P}=\{p_{1},p_{2},\ldots\}$. Define $s_{1}=p_{1}$, $s_{2}=(p_{1}p_{2})^{1/2}$, $s_{3}=(p_{1}p_{2}p_{3})^{1/3}$, and in general $s_{k}=(p_{1}\cdots p_{k})^{1/k}$ for each $k\in\mathbb{N}$. Then, for each $n\in\mathbb{N}$,
@@ 5 Ramification Indices and Residue Field Degrees
Let $\mathbb{Q}^{(n)}=\mathbb{Q}(\mathcal{A}_{n})$ be the compositum in $\overline{\mathbb{Q}}$ of all number fields of degree bounded by $n$ and let $O_{\mathbb{Q}^{(n)}}$ be the ring of integers of $\mathbb{Q}^{(n)}$. It is known that there exists a global bound on both ramification indexes and residue field degrees of valuations of $\mathbb{Q}^{(n)}$. That is, there exists $N\in\mathbb{N}$ such that if $u_{p}$ is a valuation of $\mathbb{Q}^{(n)}$ extending some $v_{p}$, $p\in\mathbb{P}$, then $e(u_{p}|v_{p})\leq N$ and $f(u_{p}|v_{p})\leq N$ [3, Proposition 4.5.3, p. 118]; we stress that $N$ is independent from $p$. So, any subset $S\subseteq O_{\mathbb{Q}^{(n)}}$ meets the conditions of Lemma 5.1, and hence $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ is nontrivial. Furthermore, subsets of $O_{\mathbb{Q}^{(n)}}$ may have unbounded degree. For an explicit example, let =$\mathbb{P}=\{p_{k}\}_{k\in\mathbb{N}}$. For each $k\in\mathbb{N}$, let $s_{k}=\sum_{i=1}^{k}\sqrt{p_{i}}$, and take $S=\{s_{k}\}_{k\in\mathbb{N}}$. Then, $S\subseteq O_{\mathbb{Q}^{(2)}}$ and for each $k$, $[\mathbb{Q}(s_{k}):\mathbb{Q}]=2^{k}$.
@@
Since overrings of Prüfer domains are Prüfer, it suffices to prove that $\textnormal{Int}_{\mathbb{Q}}(O)$ is Prüfer. By the properties of ramification indexes and residue field degree mentioned in Example 5.2, it follows that $O$ is an almost Dedekind domain with finite residue fields satisfying the double-boundedness condition of [14, Theorem 2.5]. Hence, $\textnormal{Int}(O)$ is Prüfer. By [22, Lemma 2.20], $\textnormal{Int}(O)$ is integral over $\textnormal{Int}_{\mathbb{Q}}(O)$, so $\textnormal{Int}_{\mathbb{Q}}(O)$ is Prüfer by [11, Theorem 22.4]. ∎
@@
For a fixed prime $p\in\mathbb{P}$, we consider the set $S$ of all $m$-th root of unity in $\overline{\mathbb{Z}}$ with $(p,m)=1$; we recall that $p$ is not ramified in $\mathbb{Q}(\zeta_{m})$ [16, Corollary of Theorem 26]. The set of residue field degrees of the set of prime ideals of the finite extensions generated by the elements of the set $\{p\cdot\zeta_{m}\mid m\in\mathbb{N},(p,m)=1\}$ is unbounded but $u_{p}(p\cdot\zeta_{m})>0$ for each of the relevant $m$’s. Clearly, $\frac{X}{p}\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$. The important condition here is that the residue set $S/\overline{P}$, for each prime ideal $\overline{P}\subset\overline{\mathbb{Z}}$ above $p$, is finite.
@@
    there exists some $\overline{P}\in\mathcal{P}_{p}$ such that #$\#S/\overline{P}>n$.
@@
Let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})$ be of degree $d\leq n$ and suppose that $f=\frac{g}{p}$ for some monic $g\in\mathbb{Z}_{(p)}[X]$ (see the arguments of the proof of Proposition 6.4 for the monic assumption). (i) By assumption, there exists a prime ideal $\overline{P}\in\mathcal{P}_{p}$ such that #$\#S/\overline{P}>n$, that is, there exist $s_{0},\ldots,s_{n}\in S$ such that $u_{\overline{P}}(s_{i}-s_{j})=0$ for each $0\leq i<j\leq n$. In particular, $\overline{s_{i}}=s_{i}\pmod{\overline{P}}$ are $n+1$ distinct elements of $\overline{\mathbb{F}_{p}}$. Then $g(s_{i})\equiv 0\pmod{\overline{P}}$, for $i=0,\ldots,n$, that is, $\overline{g}(\overline{s_{i}})=0$ in $\overline{\mathbb{F}_{p}}$, where $\overline{g}\in\mathbb{F}_{p}[X]$ is the reduction modulo $p$ of $g$. It follows that $\overline{g}\in\mathbb{F}_{p}[X]$ is a polynomial of degree $\leq n$ which has $n+1$ distinct roots in $\overline{\mathbb{F}_{p}}$, a contradiction.
(ii) By assumption, there exists a prime ideal $\overline{P}\in\mathcal{P}_{p}$ and $s\in S$ such that $v_{\overline{P}}(s)<\frac{1}{n}$. Let $g(X)=\prod_{i=1}^{d}(X-\alpha_{i})$ over $\overline{\mathbb{Q}}$ and let $i\in\{1,\ldots,n\}$. If $v_{\overline{P}}(\alpha_{i})$ is not 00 then it is equal to $\frac{a}{n}>0$, for some $a\in\mathbb{N}$ (not necessarily coprime with $n$) since $e(\overline{P}\cap O_{\mathbb{Q}(\alpha_{i})}\mid p)\leq[\mathbb{Q}(\alpha_{i}):\mathbb{Q}]\leq n$. If this second case occurs, then $0<v_{\overline{P}}(s)<\frac{1}{n}\leq v_{\overline{P}}(\alpha_{i})$. In either case we have:
@@
For an explicit example of a subset $S$ of $\overline{\mathbb{Z}}$ for which $S/\overline{P}$ is finite for each maximal ideal $\overline{P}\in\mathcal{P}_{p}$ for some prime $p\in\mathbb{Z}$ but {#$\{\#S/\overline{P}\mid\overline{P}\in\mathcal{P}_{p}\}$ is unbounded, let $S$ be equal to the almost Dedekind domain $D$ with finite residue fields of [12, Example 14] (see also [4, Example VI.4.18]). In that example, $D$ is the integral closure of $\mathbb{Z}_{(p)}$ in a suitable infinite algebraic extension of $\mathbb{Q}$ constructed as the union of a tower of finite algebraic extensions, so by Lemma 5.8 (see also [4, Lemma VI.4.2]) applied to that example shows that $\textnormal{Int}_{\mathbb{Q}}(D,\overline{\mathbb{Z}})=\textnormal{Int}_{\mathbb{Q}}(D)=\mathbb{Z}_{(p)}[X]$.
@@
For each $\overline{P}\in\mathcal{P}_{p}$, let $u_{\overline{P}}$ be the valuation associated to ¯$\overline{\mathbb{Z}}_{\overline{P}}$. By [19, Chapter 2, §8, Theorem 8.1], for each $\overline{P}\in\mathcal{P}_{p}$, the extension $u_{\overline{P}}$ of $v_{p}$ to $\overline{\mathbb{Q}}$ is equal to $v_{p}\circ\tau_{\overline{P}}$, for some $\mathbb{Q}$-embedding $\tau_{\overline{P}}:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{p}}$, where $v_{p}$ is the unique extension to $\overline{\mathbb{Q}_{p}}$ of the $p$-adic valuation. Given $S\subseteq\overline{\mathbb{Z}}$, we define the following subset of $\overline{\mathbb{Z}_{p}}$:
@@
()($\subseteq$)( ⊆ ) Let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}}_{\overline{P}})$ and $\tau_{\overline{P}}(s)\in\tau_{\overline{P}}(S)$, for some $s\in S$. Then $f(s)\in\overline{\mathbb{Z}}_{\overline{P}}$, so $u_{\overline{P}}(f(s))\geq 0$. Hence,
@@
()superset-of-or-equals($\supseteq$)( ⊇ ) Let $f\in\textnormal{Int}_{\mathbb{Q}}(\tau_{\overline{P}}(S),\overline{\mathbb{Z}_{p}})$ and $s\in S$. By definition we have $f(\tau_{\overline{P}}(s))\in\overline{\mathbb{Z}_{p}}$. Hence, applying $\tau_{\overline{P}}^{-1}$ to the previous relation we get $f(s)\in\overline{\mathbb{Z}}_{\overline{P}}$, and so $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}}_{\overline{P}})$.

### v2

@@ 4 Global Conditions and Examples
Recall that when $\alpha\in\overline{\mathbb{Z}}$, the index of $\alpha$ is $\iota_{\alpha}=[O_{\mathbb{Q}(\alpha)}:\mathbb{Z}[\alpha]]$. The lemma below summarizes the basic relationships among $\Psi_{p,n}$, $\iota_{\alpha}$, and integral-valued polynomials.
@@
    ([3, Theorem 3]) Modulo $p$, $\Psi_{p,n}$ is the monic least common multiple of all polynomials in $\mathbb{F}_{p}[X]$ of degree at most $n$.
@@
For (2), if $\alpha\in\mathcal{A}_{n}$ with minimal polynomial $m_{\alpha}(X)$, then $m_{\alpha}\mid\Psi_{p,n}$ modulo $p$. Hence, $\Psi_{p,n}(\alpha)\in pO_{\mathbb{Q}(\alpha)}$. To show that $\Psi_{p,n}/p\notin\textnormal{Int}_{\mathbb{Q}}(\mathcal{A}_{n+1})$, let $f\in\mathbb{F}_{p}[X]$ be monic and irreducible of degree $n+1$. Then, $f\nmid\Psi_{p,n}$ mod $p$. Next, let $F\in\mathbb{Z}[X]$ be monic, irreducible, and such that $F\equiv f$ mod $p$ (the existence of $F$ follows from Perron’s Criterion for irreducibility [28, Theorem 2.2.5]). Let $\beta\in\overline{\mathbb{Z}}$ be a root of $F$, and let $O=O_{\mathbb{Q}(\beta)}$; note that $\beta\in\mathcal{A}_{n+1}$. Suppose that $\Psi_{p,n}(\beta)/p\in\overline{\mathbb{Z}}$. Then, $\Psi_{p,n}(\beta)\in pO$, and hence $\Psi_{p,n}(\beta)\equiv 0$ mod $p$. Since the image of $\beta$ in $O/pO$ is a root of $f$, this means that $f\mid\Psi_{p,n}$ mod $p$, which is a contradiction.
@@
    Let $\sigma\in G={\rm Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and $U$ be a valuation domain of $\overline{\mathbb{Q}}$ extending $\mathbb{Z}_{(p)}$ for some $p\in\mathbb{P}$. Then $\textnormal{Int}_{\mathbb{Q}}(S,U)=\textnormal{Int}_{\mathbb{Q}}(\sigma(S),\sigma(U))$. In particular, if $S$ is $G$-invariant (i.e., $\sigma(S)=S,\forall\sigma\in G$) and $U,U^{\prime}$ are two valuation domains extending $\mathbb{Z}_{(p)}$ then $\textnormal{Int}_{\mathbb{Q}}(S,U)=\textnormal{Int}_{\mathbb{Q}}(S,U^{\prime})$ and so $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})=\textnormal{Int}_{\mathbb{Q}}(S,U)$ for any valuation domain $U$ extending $\mathbb{Z}_{(p)}$.
@@
    Assume that for each integer $n\geq 1$, there exists a finite subset $\{s_{1},\ldots,s_{m}\}\subseteq S$ such that $\mathbb{Q}(s_{i}):\mathbb{Q}]>n$ for each $1\leq i\leq m$ and $\gcd(\iota_{s_{1}},\ldots,\iota_{s_{m}})=1$. Then, $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ is trivial.
@@
Part (1) is Example 1.3. For (2), let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ of degree $d$ and let $s_{1},\ldots,s_{m}$ be the elements of $S$ with the stated properties. For each $i$, we have $\mathbb{Q}(s_{i}):\mathbb{Q}]>d$, so $\iota_{s_{i}}f\in\mathbb{Z}[X]$ by Lemma 4.2(3). The condition $\gcd(\iota_{s_{1}},\ldots,\iota_{s_{m}})=1$ then implies that $f\in\mathbb{Z}[X]$. ∎
@@
    If $m\in\mathbb{N}$ and $q$ is a prime such that $q\nmid c$, then there exists a prime $\ell>m$ such that $q\nmid\iota_{c,\ell}$.
@@
(1) This follows from Perron’s Criterion for irreducibility [28, Theorem 2.2.5].
@@
(3) Assume $p\mid c$, and use a bar to denote reduction mod $p$. Then, $\overline{f_{c,n}}(X)=X^{n}$, and $f_{c,n}(X)=X^{n}+pF_{c,n}(X)$, where $F_{c,n}(X)=\tfrac{c^{3}}{p}X^{n-1}+\tfrac{c^{2}}{p}$. Note that $F_{c,n}\in\mathbb{Z}[x]$ and $\overline{F_{c,n}}=0$. By Theorem 4.8, $p\mid\iota_{p,n}$.
@@
(5) We have $f_{c,n}^{\prime}(X)=X^{n-2}((n-1)c^{3}+nX)$. By the Dirichlet’s theorem on arithmetic progressions, there exists a prime $\ell$ of the form $qk+1$, $\ell>m$; in particular, $v_{q}(\ell-1)>0$ and so, modulo $q$, the only root of $f_{c,\ell}^{\prime}(X)$ is 0 (note that $v_{q}(\ell)=0$) which is not a root of $f_{c,\ell}$ modulo $q$ because we are assuming $v_{q}(c)=0$. Hence, $f_{c,\ell}$ is separable modulo $q$ so that $q\nmid\iota_{c,\ell}$ by Theorem 4.8. ∎
@@
By Lemma 4.10(2), $u_{p}(s_{i})=2/q_{i}$ for each $i$, so $\{u_{p}(s_{i})\}_{i\in\mathbb{N}}$ is a strictly decreasing sequence with limit 0. By Lemma 4.6, $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})=\mathbb{Z}_{(p)}[X]$.
Let now $q$ be a prime different from $p$. Given $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$, by Lemma 4.10(5), there exists a prime $q_{i}$ such that $\mathbb{Q}(s_{i}):\mathbb{Q}]=q_{i}>\deg(f)$ and $q\nmid\iota_{s_{i}}$. By Lemma 4.2(3), $\iota_{s_{i}}\cdot f\in\mathbb{Z}[X]$ so that $f$ belongs to $\mathbb{Z}_{(q)}[X]$. This argument shows that $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(q)}})=\mathbb{Z}_{(q)}[X]$ for every prime $q\not=p$.
@@
We recall that the prime $p$ is totally ramified in $\mathbb{Q}(\zeta_{p^{k}})$ for every $k\in\mathbb{N}$ (see for example [16, Theorem 26]), and $u_{p}(1-\zeta_{p^{k}})=(p^{k-1}(p-1))^{-1}$ by [30, Chapter 2, p. 9]. Moreover, if $q$ is a prime different from $p$ then $u_{q}(1-\zeta_{p^{k}})=0$. These calculations prove the claims about $S$.
@@
Given primes $p<q$, we have $\zeta_{q}-\zeta_{p}=\zeta_{q}(1-\zeta_{pq}^{q-p})$, and $\zeta_{pq}^{q-p}$ is a primitive $pq$-th root of unity because qpq-p is coprime to $pq$. So, for any prime $l$,
@@
By [30, Proposition 2.8], $1-\zeta_{pq}$ is a unit of $\mathbb{Z}[\zeta_{pq}]$. Hence, $u_{l}(\zeta_{q}-\zeta_{p})=0$, and so for every prime $l$, $S$ is pseudo-stationary with respect to $u_{l}$ with breadth ideal $U_{l}$.
@@
Let $\mathbb{P}=\{p_{1},p_{2},\ldots\}$. Define $s_{1}=p_{1}$, $s_{2}=(p_{1}p_{2})^{1/2}$, $s_{3}=(p_{1}p_{2}p_{3})^{1/3}$, and in general $s_{k}=(p_{1}\cdots p_{k})^{1/k}$ for each $k\in\mathbb{N}$. Then, for each $n\in\mathbb{N}$,
@@ 5 Ramification Indices and Residue Field Degrees
Let $\mathbb{Q}^{(n)}=\mathbb{Q}(\mathcal{A}_{n})$ be the compositum in $\overline{\mathbb{Q}}$ of all number fields of degree bounded by $n$ and let $O_{\mathbb{Q}^{(n)}}$ be the ring of integers of $\mathbb{Q}^{(n)}$. It is known that there exists a global bound on both ramification indexes and residue field degrees of valuations of $\mathbb{Q}^{(n)}$. That is, there exists $N\in\mathbb{N}$ such that if $u_{p}$ is a valuation of $\mathbb{Q}^{(n)}$ extending some $v_{p}$, $p\in\mathbb{P}$, then $e(u_{p}|v_{p})\leq N$ and $f(u_{p}|v_{p})\leq N$ [2, Proposition 4.5.3, p. 118]; we stress that $N$ is independent from $p$. So, any subset $S\subseteq O_{\mathbb{Q}^{(n)}}$ meets the conditions of Lemma 5.1, and hence $\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$ is nontrivial. Furthermore, subsets of $O_{\mathbb{Q}^{(n)}}$ may have unbounded degree. For an explicit example, let $\mathbb{P}=\{p_{k}\}_{k\in\mathbb{N}}$. For each $k\in\mathbb{N}$, let $s_{k}=\sum_{i=1}^{k}\sqrt{p_{i}}$, and take $S=\{s_{k}\}_{k\in\mathbb{N}}$. Then, $S\subseteq O_{\mathbb{Q}^{(2)}}$ and for each $k$, $\mathbb{Q}(s_{k}):\mathbb{Q}]=2^{k}$.
@@
Since overrings of Prüfer domains are Prüfer, it suffices to prove that $\textnormal{Int}_{\mathbb{Q}}(O)$ is Prüfer. By the properties of ramification indexes and residue field degree mentioned in Example 5.2, it follows that $O$ is an almost Dedekind domain with finite residue fields satisfying the double-boundedness condition of [14, Theorem 2.5]. Hence, $\textnormal{Int}(O)$ is Prüfer. By [23, Lemma 2.20], $\textnormal{Int}(O)$ is integral over $\textnormal{Int}_{\mathbb{Q}}(O)$, so $\textnormal{Int}_{\mathbb{Q}}(O)$ is Prüfer by [11, Theorem 22.4]. ∎
@@
For a fixed prime $p\in\mathbb{P}$, we consider the $m^{\text{th}}$ roots of unity in $\overline{\mathbb{Z}}$ with $(p,m)=1$; we recall that $p$ is not ramified in $\mathbb{Q}(\zeta_{m})$ [16, Corollary of Theorem 26]. The set of residue field degrees of the set of prime ideals of the finite extensions generated by the elements of the set $S=\{p\cdot\zeta_{m}\mid m\in\mathbb{N},(p,m)=1\}$ is unbounded but $u_{p}(p\cdot\zeta_{m})>0$ for each of the relevant $m$’s. Clearly, $\frac{X}{p}\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}})$. The important condition here is that the residue set $S/\overline{P}$, for each prime ideal $\overline{P}\subset\overline{\mathbb{Z}}$ above $p$, is finite.
@@
    there exists some $\overline{P}\in\mathcal{P}_{p}$ such that $S/\overline{P}>n$.
@@
Let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}_{(p)}})\setminus\mathbb{Z}_{(p)}[X]$ be of degree $d\leq n$; without loss of generality, we may assume that $f=\frac{g}{p}$ for some monic $g\in\mathbb{Z}_{(p)}[X]$ (see the arguments of the proof of Proposition 6.4 for the monic assumption; we may multiply $f$ by a suitable power of $p$ in order to get $p$ at the denominator).
(i) By assumption, there exists a prime ideal $\overline{P}\in\mathcal{P}_{p}$ such that $S/\overline{P}>n$, that is, there exist $s_{0},\ldots,s_{n}\in S$ such that $u_{\overline{P}}(s_{i}-s_{j})=0$ for each $0\leq i<j\leq n$. In particular, $\overline{s_{i}}=s_{i}\pmod{\overline{P}}$ are $n+1$ distinct elements of $\overline{\mathbb{F}_{p}}$. Then $g(s_{i})\equiv 0\pmod{\overline{P}}$, for $i=0,\ldots,n$, that is, $\overline{g}(\overline{s_{i}})=0$ in $\overline{\mathbb{F}_{p}}$, where $\overline{g}\in\mathbb{F}_{p}[X]$ is the reduction modulo $p$ of $g$. It follows that $\overline{g}\in\mathbb{F}_{p}[X]$ is a polynomial of degree $\leq n$ which has $n+1$ distinct roots in $\overline{\mathbb{F}_{p}}$, a contradiction.
(ii) By assumption, there exists a prime ideal $\overline{P}\in\mathcal{P}_{p}$ and $s\in S$ such that $v_{\overline{P}}(s)<\frac{1}{n}$. Let $g(X)=\prod_{i=1}^{d}(X-\alpha_{i})$ over $\overline{\mathbb{Q}}$, where each $\alpha_{i}\in\mathcal{A}_{n}$. Fix $i\in\{1,\ldots,n\}$. If $v_{\overline{P}}(\alpha_{i})$ is not 0 then it is equal to $\frac{a}{n}>0$, for some $a\in\mathbb{N}$ (not necessarily coprime with $n$) since $e(\overline{P}\cap O_{\mathbb{Q}(\alpha_{i})}\mid p)\leq[\mathbb{Q}(\alpha_{i}):\mathbb{Q}]\leq n$. If this second case occurs, then $0<v_{\overline{P}}(s)<\frac{1}{n}\leq v_{\overline{P}}(\alpha_{i})$. In either case we have:
@@
For an explicit example of a subset $S$ of $\overline{\mathbb{Z}}$ for which $S/\overline{P}$ is finite for each maximal ideal $\overline{P}\in\mathcal{P}_{p}$ for some prime $p\in\mathbb{Z}$ but $\{\#S/\overline{P}\mid\overline{P}\in\mathcal{P}_{p}\}$ is unbounded, let $S$ be equal to the almost Dedekind domain $D$ with finite residue fields of [12, Example 14] (see also [4, Example VI.4.18]). In that example, $D$ is the integral closure of $\mathbb{Z}_{(p)}$ in a suitable infinite algebraic extension of $\mathbb{Q}$ constructed as the union of a tower of finite algebraic extensions, so by Lemma 5.8 (see also [4, Lemma VI.4.2]) applied to that example shows that $\textnormal{Int}_{\mathbb{Q}}(D,\overline{\mathbb{Z}})=\textnormal{Int}_{\mathbb{Q}}(D)=\mathbb{Z}_{(p)}[X]$.
@@
For each $\overline{P}\in\mathcal{P}_{p}$, let $u_{\overline{P}}$ be the valuation associated to $\overline{\mathbb{Z}}_{\overline{P}}$. By [19, Chapter 2, §8, Theorem 8.1], for each $\overline{P}\in\mathcal{P}_{p}$, the extension $u_{\overline{P}}$ of $v_{p}$ to $\overline{\mathbb{Q}}$ is equal to $v_{p}\circ\tau_{\overline{P}}$, for some $\mathbb{Q}$-embedding $\tau_{\overline{P}}:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{p}}$, where $v_{p}$ is the unique extension to $\overline{\mathbb{Q}_{p}}$ of the $p$-adic valuation. Given $S\subseteq\overline{\mathbb{Z}}$, we define the following subset of $\overline{\mathbb{Z}_{p}}$:
@@
$(\subseteq)$ Let $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}}_{\overline{P}})$ and $\tau_{\overline{P}}(s)\in\tau_{\overline{P}}(S)$, for some $s\in S$. Then $f(s)\in\overline{\mathbb{Z}}_{\overline{P}}$, so $u_{\overline{P}}(f(s))\geq 0$. Hence,
@@
$(\supseteq)$ Let $f\in\textnormal{Int}_{\mathbb{Q}}(\tau_{\overline{P}}(S),\overline{\mathbb{Z}_{p}})$ and $s\in S$. By definition we have $f(\tau_{\overline{P}}(s))\in\overline{\mathbb{Z}_{p}}$. Hence, applying $\tau_{\overline{P}}^{-1}$ to the previous relation we get $f(s)\in\overline{\mathbb{Z}}_{\overline{P}}$, and so $f\in\textnormal{Int}_{\mathbb{Q}}(S,\overline{\mathbb{Z}}_{\overline{P}})$.